OpenStudy (anonymous):

use the quadratic formula to slove the equation -4x^2-3x+2=0

OpenStudy (anonymous):

\[-4x^2-3x+2 =0 \rightarrow 4x^2+3x-2 =0 \rightarrow ax^2+bx+c =0\] a= 4, b= 3 & c= -2 Now quadratic formula is given as; \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} =\frac{-3 \pm \sqrt{3^2-4\times 4 (-2)}}{2 \times 4} \] \[=\frac{-3 \pm \sqrt{9+32}}{8} =\frac{-3 \pm \sqrt{41}}{8} \] i.e. either \[x=\frac{-3 + \sqrt{41}}{8} --- or --x=\frac{-3- \sqrt{41}}{8}\] are the required solutions of the given quadratic equation. @NOOB_ZERO456

OpenStudy (anonymous):

this is a multiple choice question \[-\frac{ 3 }{ 8 }\pm \frac{ \sqrt{41} }{ 8 }\] \[-\frac{ 8 }{ 3}\pm \frac{ \sqrt{82} }{ 8 }\] \[-\frac{ 3 }{ 8 }\pm \frac{ \sqrt{20} }{2}\] \[-\frac{ 3 }{ 4 }\pm \frac{ \sqrt{41} }{ 4 }\] i don't understand this and would appreciate some help explaining this to me

OpenStudy (anonymous):

First option is the correct answer @NOOB_ZERO456

OpenStudy (anonymous):

thanks you

OpenStudy (anonymous):

\[ \huge x=\frac{−3± \sqrt{41}}{8} =\frac{−3}{8}± \frac{ \sqrt{41}}{8}\] @NOOB_ZERO456 medal please